OpenStudy (anonymous):

Use a related equation that has the variable on one side. Then simplify the other side. m over 4= 16 A. m = 64 B. m = 4 C. m = 12 D. m = 20

OpenStudy (nubeer):

multiply both sides by 4. m = 4*16

OpenStudy (kirbykirby):

\[ \frac{m}{4}=16\] What can you do to both sides of the equation such that you only get "m" on the left side?

OpenStudy (anonymous):

idk

OpenStudy (kirbykirby):

As nubeer said, if you multiply by 4 on both sides: you get \[ 4\frac{m}{4}=4(16)\\ m = 4(16)\]

OpenStudy (anonymous):

a

OpenStudy (kirbykirby):

yes

OpenStudy (anonymous):

ok yhinks